Assume that one adult was smart phones are randomly selected 64% use them in meetings are classes if not adult smart phone users are randomly selected find the probability that exactly 2 of them use their smart phones in meetings or classes


1
Expert's answer
2021-10-18T16:27:17-0400

p = 64 % = 0.64

Let 7 adult smartphone users are randomly selected,

n = 7


Then, the probability that exactly 2 of them use their smartphones in meetings or classes is


"P(X=x) = \dfrac{n!}{x!(n-x)!} \times p^x \times (1-p)^{n-x} \\\ \\

p(X=2) = \dfrac{7!}{2!(7-2)!} \times 0.64^2 \times (1-0.64)^{7-2} \\\ \\

= 21 \times 0.64^2 \times 0.36^5 \\

= 21\times 0.4096\times 0.00604\\=0.052"